OpenStudy (anonymous):

Find the distance between these points. W(-6, -8), X(6, 8) 20 10 √8

OpenStudy (anonymous):

Ok. This is simple. Do you know the distance formula?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

\[\sqrt{(x2-x1}^2 + (y2-y1)^2\] That's the formula!

OpenStudy (anonymous):

what do i do from here how do I get the anser

OpenStudy (anonymous):

Just plug it in. The points are (x1,y1) and (x2,y2)

OpenStudy (anonymous):

\[\sqrt{(8+6)^2 + (6+8)^2}\]

OpenStudy (anonymous):

Now solve.

OpenStudy (anonymous):

14+14 = 28 ???????

OpenStudy (anonymous):

No. 14^2 + 14^2

OpenStudy (anonymous):

That 14^2=14x14=196

OpenStudy (anonymous):

So 196+196

OpenStudy (anonymous):

So 392

OpenStudy (anonymous):

Square root of 392 is 19.80 Hope you understand :D What grade are you in?

OpenStudy (anonymous):

Im in 7 but how is the answer 19.80 if the answers should be one of these 20 10 √8

OpenStudy (anonymous):

Round. So the answer is 20. The first one :D

OpenStudy (anonymous):

Medals are appreciated :P

OpenStudy (anonymous):

oh okay thanks I will give you a medal :) for helping me

OpenStudy (anonymous):

Do you understand though? And your welcome! Btw I am in 8th and I am taking Geometry! So i'm doing this in my school now!